Temperature of the Fahrenheit scale is 54 F, and what is the fahrenhiet into celsius scale?

Grade:11

3 Answers

Rinkoo Gupta
10 years ago
C/5=(F-32)/8

C/5=(54-32)/9=22/9
=>C=110/9
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Rinkoo Gupta
10 years ago
C/5=(F-32)/9

C/5=(54-32)/9
=>C/5=22/9
=>C=110/9
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
sivakrishna
10 years ago
C/5=(F-32)/9 =>C/5=(54-32)/9 =>C/5=22/9 =>C=22/9*5 =>C=110/9 =>C=12.12 degree centigrade